The average of marks obtained by 120 candidates was 35. If the average of marks of passed candidates was 39 and that of failed candidates was 15, the number of candidates who passed examination is
Answer by ankor@dixie-net.com(22740)

The average of marks obtained by 120 candidates was 35.
If the average marks of passed candidates was 39 and that of failed candidates was 15, the number of candidates who passed examination is
:
let p = no. students who passed
there were 120 students, therefore
(120-p) = no. of students that failed
:
Find the total no. of marks of the 120 students
120 * 35 = 4200
therefore
39p + 15(120-p) = 4200
39p + 1800 - 15p = 4200
39p - 15p = 4200 - 1800
24p = 2400
p = 2400/24
p = 100 students passed.
;
:
Check, find the sum of the total marks of each group
100*39 = 3900
20 *15 = 3000
--------------
total: 4200